Factor: 4p^2 - 8p - 60
Answer by nabla(475)

First, all terms have a 4 in common, factor that out first.
4p^2 - 8p - 60=4(p^2-2p-15)
Now, we need two numbers that multiply to get -15 but ADD to get -2. These would be -5 and 3.
So,
4(p^2-2p-15)=4(p-5)(p+3)